OpenStudy (anonymous):

gary ozawa gave you the following directions to get form school to his home. go six blocks east, then 15 blocks north, then four more blocks east, then 1 block south, and finally 1 block west. by air to his home, about how many kilometers is it, if 5 blocks is about 1 kilometer?

Directrix (directrix):

Walking the following number of blocks and directions 6 E, 15 N, 4 E, 1 S, and 1 W will take a person to the same location as if the person had walked 9 blocks E and 14 blocks N. These lengths are those of the legs of a right triangle having hypotenuse the distance as the "crow flies" (air distance) from the walker's starting to ending point. 9^2 + 14^2 = d^2 81 + 196 = d^2 277 = d^2 16.6433 = d The distance is in blocks. Given that one kilometer is approximately 5 blocks, 16.6433 blocks is roughly equivalent to 3.33 km.

OpenStudy (anonymous):

oh, so you just simplify the numbers then?

OpenStudy (anonymous):

I think you mean 10 blocks east, 14 north

Directrix (directrix):

I meant 9 E and 14 N. That comes from (10 E - 1 W ) and (15 N - 1 S). The route will not be the same as the one walked by the Gary O but a single right triangle will be created. I'll see if I can draw a diagram.

Directrix (directrix):